A tightly stretched string with fixed end points x = 0 and x = l is initially in a position

given by y = y0 sin3(πx/l) . It is released from rest from the initial position. Find the

displacement y(x, t) .


1
Expert's answer
2021-06-07T13:06:56-0400

Wave equation:

"y_{xx}=y_{tt}\/c^2"

Separate the variables:

"y(x,t)=X(x)Y(t)"

"\frac{X''}{X}=\frac{Y''}{c^2Y}=-k^2"


"X''+k^2X=0"

"X(x)=c_1sin(kx)+c_2cos(kx)"


"Y''+c^2k^2Y=0"

"Y(t)=c_3sin(ckt)+c_4cos(ckt)"


"y(x,t)=(c_1sin(kx)+c_2cos(kx))(c_3sin(ckt)+c_4cos(ckt))"


The boundary conditions:

"y(0,t)=0" . fixed node

"y(l,t)=0" . fixed node

"y_t(x,0)=0" , string is only released at t=0

"y(x,0)=y_0sin^3(\pi x/l)" , initial shape


For the 1st condition:

"c_2(c_3sin(ckt)+c_4cos(ckt))=0"

For the 2nd condition:

"c_1sin(kl)(c_3sin(ckt)+cos(ckt))=0"

"k=n\pi/l" , "n=1,2,..."


Now, the solution:

"y(x,t)=c_1sin(n\pi x/l)(c_3sin(n\pi ct/l)+c_4cos(n\pi ct/l))"


For the 3rd condition:

"c_1sin(n\pi x/l)(c_3n\pi c/l)=0\implies c_3=0"


"y(x,t)=c_1c_4\displaystyle{\sum^{\infin}_1b_nsin(n\pi x/l)cos(n\pi ct/l)}"


For the 4th condition:

"y(x,0)=y_0sin^3(\pi x/l)=y_0(\frac{3}{4}sin(\pi x/l)-\frac{1}{4}sin(3\pi x/l))="

"=\displaystyle{\sum^{\infin}_1b_nsin(n\pi x/l)}"

Comparing coefficients, we get:

"b_1=3y_0\/4"

"b_2=b_4=0"

"b_3=-y_0\/4"


The final solution:

"y(x,t)=\frac{3}{4}y_0(\pi x/l)cos(\pi ct/l)-\frac{1}{4}y_0(3\pi x/l)cos(3\pi ct/l)"